What is the pH at which a buffer composed of NH3 and NH4+ would be most effective at resisting pH change? Round your answer to 2 decimal places.
Answer #1

The buffer is most effective when pH is equal to pKa

Kb of NH3 is 1.8*10^-5

lets find Ka of NH4+ and then its pKa

use:

Ka = Kw/Kb

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Ka = (1.0*10^-14)/Kb

Ka = (1.0*10^-14)/1.8*10^-5

Ka = 5.556*10^-10

use:

pKa = -log Ka

= -log (5.556*10^-10)

= 9.2553

Answer: 9.26